[Shuta Tanaka, Ide, Obayashi (Aogaku), Yuzurihara]
We performed the lockloss investigation for the lockloss which happened between 2025/06/11~06/17 (first week). The previous lockloss investigation was posted in klog34687. During this period, there were 36 lockloss from the OBSERVATION state. There was no new phenomena. The plots are listed in DAC wiki.
ETMY MN oplev glitch
As reported in klog34259, in the begining after joining O4c, the ETMY MN oplev glitch oftern happened. In the previous report, we found the glitch behavior around 11 lockloss. We carefully checked the data to judge whether this glitch was the cause of the lockloss.
When there was a sudden fluctuation in K1:VIS-ETMY_MN_COILOUTF_V3_OUTPUT above 4000, we judged this glitch was the cause of the lockloss. This threshold was determined empirically. We checked the behavior of other channels, but did not find nothing that seemed to be related. As a result, 4 lockloss (i=1, 2, 4, 5)was caused by this glitch.
Details was sunnaruzed in JGWdoc (in Japanese).
Quick saturation of OMC DCPD
For 2 lockloss (i=16, 34), the quick saturation of OMC DCPD happened. This phenomena looks similar with the lockloss reported in klog34527. Compared with the usual saturation of OMC DCPD, the quickness (or the time scale) of the saturation is different.
Slowly saturation of OMC DCPD
For 22 lockloss, the slowly saturation of OMC DCPD can be seen. The oscillation just before the saturation was 20~40 Hz. Three of them have the ISC_WC_L label.
There was no lockloss with 18 Hz oscillation.
Earthquake
For 3 lockloss, the EQ label was asgned.
Excess of 1~10Hz seismic motion
For 4 lockloss (Three of them has also EQ label), 1Hz label was assigned. In addition, for 1 lockloss (i=27), we found the excess of the 1Hz seismic motion.
Alignment upward the AS port
There was no lockloss.
Lockloss by the human request
For 3 lockloss, the REQ label was asigned.